# Visual Instruction Pretraining for Domain-Specific Foundation Models

### v2

@@ Abstract
Modern computer vision is converging on a closed loop in which perception, reasoning and generation mutually reinforce each other. However, this loop remains incomplete: the top-down influence of high-level reasoning on the foundational learning of low-level perceptual features is not yet underexplored. This paper addresses this gap by proposing a new paradigm for pretraining foundation models in downstream domains. We introduce Visual insTruction Pretraining (ViTP), a novel approach that directly leverages reasoning to enhance perception. ViTP embeds a Vision Transformer (ViT) backbone within a Vision-Language Model and pretrains it end-to-end using a rich corpus of visual instruction data curated from target downstream domains. ViTP is powered by our proposed Visual Robustness Learning (VRL), which compels the ViT to learn robust and domain-relevant features from a sparse set of visual tokens. Extensive experiments on 16 challenging remote sensing and medical imaging benchmarks demonstrate that ViTP establishes new state-of-the-art performance across a diverse range of downstream tasks. The code is available at https://github.com/zcablii/ViTP.
@@ 1 Introduction
Figure 1: The synergistic relationship between perception, generation, and reasoning in modern CV. Our proposed ViTP forges a novel link from high-level reasoning to low-level perception, a previously underexplored connection.
Recent years have witnessed a leap forward in computer vision (CV), largely catalyzed by the Transformer architecture [132]. This architecture has emerged as a unifying framework across the three cornerstone areas of CV: perception, generation, and reasoning. A synergistic relationship has become evident, where advancements in one area often precipitate progress in the others, as illustrated in Figure 1. For instance, generative pretraining methods like Masked Image Modeling (MIM) [50, 6] have significantly boosted the performance of downstream perception tasks. Conversely, perception-centric features from models like DINOv2 [108] have been leveraged to guide and accelerate generative diffusion model training [145]. Such interplay is also demonstrated by the modern Vision-Language Models (VLMs) [59, 3, 21], ImageGen-CoT [81] and LLaDa [106]. However, a crucial link remains underexplored: the top-down influence of high-level reasoning on the foundational learning of low-level perceptual features. This raises a compelling question: can abstract reasoning, which often demands highly discriminative visual features, be harnessed to guide a perception model toward learning more precise and efficient representations?
@@
Figure 2: Comparison of pretraining paradigms for ViT foundation models. ViTP employs an instruction-following objective to directly instil domain-specific perception capabilities into the vision backbone.
Answering this question is particularly critical for pretraining powerful foundation models in specialized domains like remote sensing and medical imaging, where prevailing paradigms face formidable challenges. Prevailing pretraining paradigms include supervised classification, MIM [50, 6, 112], image-image contrastive learning [49, 108, 41], and image-text contrastive learning [110, 88, 171]. Each approach, however, presents notable drawbacks. Supervised pretraining risks overfitting to narrow label distributions, thereby limiting generalization. MIM, by focusing on pixel-level reconstruction, may inadvertently neglect the fine-grained details of small but critical objects, a common scenario in remote sensing and medical scans. Image-image contrastive methods, while effective, are often notoriously difficult to optimize and demand substantial computational resources. These paradigms are all fundamentally “bottom-up”: they operate on the premise that robust low-level perception is a prerequisite for high-level reasoning [93, 84]. While this hierarchy is well-established, the reciprocal pathway (i.e. whether high-level reasoning can directly enhance low-level feature learning) remains largely unexplored.
@@
Image-to-text contrastive learning, exemplified by CLIP [110], represents an initial step toward reasoning-guided pretraining. However, these methods primarily aim for general-purpose image-text alignment rather than optimizing a vision backbone for specialized, fine-grained downstream tasks. Consequently, the global, image-level features learned by models like CLIP’s ViT often prove suboptimal for dense prediction tasks such as semantic segmentation or object detection, which demand precise, pixel-level understanding [79]. This limitation underscores that merely aligning global image and text representations is an insufficient strategy for forging a powerful, domain-specific perception backbone.
@@
To bridge this gap, we introduce Visual insTruction Pretraining (ViTP), a novel, top-down pretraining paradigm that directly integrates high-level, instruction-based reasoning into the perceptual feature learning process of a ViT backbone. As illustrated in Figure 2, ViTP diverges from prior methods by embedding the ViT within a larger Vision-Language Model (VLM) and pretraining it via a visual instruction-following objective. We construct our training data with image-text pairs derived directly from the target downstream tasks. During training, image tokens from the ViT and text tokens from a user query are fed jointly into a Large Language Model (LLM) [89]. The LLM’s generated response serves as the supervisory signal, compelling the ViT backbone to learn the complex data distributions of the downstream domain in an end-to-end fashion. This process endows the ViT with highly relevant, domain-specific perceptual capabilities. Furthermore, we introduce Visual Robustness Learning (VRL) by randomly dropping a large fraction of the ViT’s output image tokens before they are passed to the LLM. This constraint implicitly forces the ViT’s attention mechanism to encode more comprehensive and robust information within each of the remaining tokens, thereby enhancing the robustness and semantic richness of the learned visual features.
@@
  • We introduce ViTP, a top-down pretraining paradigm that leverages instruction following objectives to imbue a ViT backbone with high-level semantic perception tailored for specific downstream domains.
  • We propose Visual Robustness Learning, a regularization method that encourages the ViT to learn more comprehensive and robust feature representations by operating on a sparse set of visual tokens.
@@ 2.1 Perception Foundation Model
2.1.1 General Domains
Early approaches predominantly relied on supervised pretraining over large-scale labeled datasets such as ImageNet [47, 34]. While effective in learning semantically meaningful representations, these methods risk overfitting to the label space of the pretraining task, thereby limiting generalization to diverse downstream applications. Consequently, the field has increasingly shifted toward unsupervised learning paradigms that leverage unlabeled data. Among these, contrastive learning and masked image modeling (MIM) emerged as two dominant strategies. Contrastive learning methods, including MoCo [49], BYOL [39], and DINOv2 [108], learn discriminative representations by minimizing the distance between positive pairs (e.g., augmented views of the same image) while maximizing the distance between negative pairs in the embedding space. These methods emphasize semantic invariance and shown strong generalization capabilities across tasks. On the other hand, MIM-based approaches such as MAE [50], SimMIM [150] and UM-MAE [71] adopt a generative pretext task: a large portion of the input image is masked, and the model is trained to reconstruct the missing content. This encourages the model to learn holistic, contextually rich visual features by capturing spatial dependencies and fine-grained details.
While both paradigms achieved impressive results on natural images, their direct transfer to specialized domains such as remote sensing or medical imagery often results in suboptimal performance [27, 88, 42]. This underscores the necessity for domain-specific adaptations in pretraining strategies.
2.1.2 Remote sensing
2.1.3 Medical Images
2.2 Continual Pretraining
@@ 2.3 Visual Instruction Tuning
The rise of powerful VLMs such as Gemini [127], InternVL [20], and Qwen-VL [3] are largely driven by visual instruction tuning [89]. This paradigm aims to align visual representations with a pretrained Large Language Model (LLM) by projecting image features into the LLM’s embedding space. The model is then finetuned on a corpus of image-text instruction pairs, enabling the LLM to interpret visual inputs and respond to textual instructions. In this setup, the vision encoder typically remains frozen, serving solely as a feature extractor, while all learning occurs in the LLM and the projection module.
While this paradigm effectively leverages perception to enhance reasoning, the reverse direction, i.e. using reasoning to improve perception, remains largely underexplored. Our proposed ViTP addresses this gap by inverting the traditional visual instruction tuning pipeline. Instead of using images to tune an LLM, we utilize an LLM to guide the pretraining of the vision encoder.
@@ 3 Method
Figure 4: A conceptual illustration of the ViTP framework. A ViT backbone is embedded within a large VLM and then pretrained with domain-specific instruction following objective and Visual Robustness Learning (VRL). This process instils high-level semantic understanding into the ViT. The resulting weights are then used to initialize models for various downstream perception tasks.
In this section, we present our visual instruction pretraining (ViTP) framework, a novel paradigm designed to pretrain a ViT backbone. As illustrated in Figure 2, ViTP fundamentally differs from traditional pretraining by leveraging the reasoning capabilities of modern VLMs. The entire process is driven by a “visual instruction following objective”, where the ViT learns to extract features that help an LLM answer questions about an image. To tailor the model for specific downstream applications, we outline a “data recipe” for curating a domain-specific pretraining dataset. During pretraining, we employ “Visual Robustness Learning (VRL)”, a regularization technique that drops image tokens to enhance the semantic robustness of the learned representations. The final pretrained ViT serves as a powerful foundation model, readily adaptable to a variety of downstream tasks.
@@ 3.1 Visual Instruction Following Objective
The central hypothesis of ViTP is that a ViT can learn more potent and relevant features if its training is guided by high-level reasoning. To achieve this, we frame the pretraining as a visual instruction-following task. The framework, depicted in Figure 4, processes a domain-specific image through a ViT encoder to produce a sequence of image tokens. These tokens are projected into the LLM’s embedding space and concatenated with the tokenized text of an instruction. A Large Language Model (LLM) then processes this combined sequence to generate a response. The entire model is trained end-to-end, allowing the supervisory signal from the LLM’s response to directly optimize the ViT’s feature extraction process.
Our approach employs a continual pretraining strategy [43], starting with a well-trained, general-purpose VLM. We then continue its training on our curated domain-specific datasets. This offers two primary advantages: 1) It leverages the VLM’s vast pre-existing knowledge of general visual and linguistic patterns, providing a robust initialization. 2) This strong starting point significantly accelerates convergence during domain-specific pretraining, enhancing computational efficiency. Let the raw dataset be $\mathcal{D}_{\text{raw}}=\{(I,Q,R)\}$, where $I$ is an image, $Q$ is a text query (instruction), and $R$ is the ground-truth text response. This is processed into the final training set $\mathcal{D}=\{(x_{i},x_{t},y^{\star})\}$, where $x_{i}$, $x_{t}$, and $y^{\star}$ represent the processed image tokens, text tokens, and target response tokens, respectively.
@@ 3.1.3 LLM-based Supervision
The combined sequence $S_{llm}$ is processed by the LLM, which acts as a reasoning engine to interpret the visual information from $x_{i}$ in the context of the instruction $x_{t}$. The model then auto-regressively generates an output sequence $O$. The entire system is optimized by minimizing the discrepancy between the generated output $O$ and the ground-truth target response $y^{\star}$. For instance, if the instruction asks to identify an object, $y^{\star}$ could be a textual description including its location in a structured format. During pretraining, the gradients from the output loss propagate back through the entire model, including the projection layer and the ViT backbone. We allow the weights of the ViT, the projection layer, and the LLM to be trainable. The optimization follows a standard supervised finetuning (SFT) objective, which minimizes the negative log-likelihood of the target sequence:
@@ 3.3 Pretraining Dataset Recipe
    Task Capability Alignment: The instruction-following tasks in the pretraining data should foster abilities required downstream. For instance, for downstream object detection tasks, including visual grounding and fine-grained VQA during pretraining empowers the backbone with localization and spatial reasoning capabilities.
@@ 3.4 Downstream Finetuning
Once ViTP is complete, the pretrained ViT backbone is then extracted to serve as a powerful backbone network. For downstream applications, this pretrained ViT is integrated into a standard ViT-Adapter [18] and combined with task-specific heads for detection, segmentation, or other tasks. The entire model is then finetuned on the target downstream dataset. This transfer learning approach leverages the rich, instruction-aware representations learned during ViTP, enabling faster adaptation and superior performance on specialized tasks.
Table 1: Overview of downstream task datasets for remote sensing, including object detection (horizontal and oriented), segmentation (semantic and instance), and change detection.
Task Dataset Modal Box Format Images Classes
  Object Detection DIOR [63] RGB Hori. Box 23,463 20
DIOR-R [25] RGB Ori. Box 23,463 20
DOTA-v2.0 [146] RGB Ori. Box 11,268 18
SARDet-100K [77] SAR Hori. Box 116,598 6
SSDD [167] SAR Hori. Box 1,160 1
RSAR [170] SAR Ori. Box 95,842 6
Semantic and Instance Segmentation iSAID [143] RGB Mask 2,806 15
LoveDA [139] RGB Mask 5,987 7
UAVid [99] RGB Mask 5,510 8
SSDD [167] SAR Polygons 1,160 1
Change Detection SVCD [62] RGB Mask 16,000 2
WHU [57] RGB Mask 11,456 2
LEVIR-CD [11] RGB Mask 10,192 2
S2Looking [115] RGB Mask 5,000 2
Medical Segmentation AMOS2022 [58] CT Mask 19,310 14
BraTS2021 [30] MRI Mask 76,467 3
CovidQUEx [124] X-ray Mask 5,826 2
3.5 Implementation Details
Training a VLM from scratch conventionally requires a multi-stage pipeline: vision–language contrastive pretraining, projector alignment, and large-scale instruction tuning. To bypass the prohibitive cost of these initial phases, we instead bootstrap from a publicly available, high-capacity VLM. Specifically, we initialise our weights from InternVL-2.5 [19], which uses a custmized ViT-Large backbone [20] and a Qwen2 [2] language model. All pretraining are performed on 8× NVIDIA A40 (48 GB) GPUs with a global batch size of 128, and all downstream task finetunings are conducted on $8\times$ NVIDIA RTX3090 (24 GB) GPUs. For ViTP pretraining, we use the AdamW optimizer with a learning rate of 2e-5 and a cosine decay schedule over 8,000 training steps. To accelerate both pre-training and downstream finetuning, Flash-Attention [29] is integrated into every ViT self-attention layer, yielding extra throughput gain without altering convergence behaviour.

### v3

@@ Abstract
Human visual perception is the product of a synergistic interplay between bottom-up perception processing and top-down understanding-driven modulation. While modern computer vision has excelled at creating powerful models that mirror the bottom-up, hierarchical nature of perception, the reciprocal top-down influence of high-level understanding on the foundational learning of low-level perception features remains critically underexplored. This paper addresses this gap by proposing a new paradigm for pretraining foundation models, inspired by the top-down mechanisms of human vision. We introduce Visual insTruction Pretraining (ViTP), a novel approach that directly leverages understanding to enhance perception. ViTP embeds a Vision Transformer backbone within a Vision-Language Model and pretrains it end-to-end using a rich corpus of visual instruction data curated from target downstream domains. ViTP is powered by our proposed Visual Robustness Learning (VRL), which compels the Vision Transformer to learn robust and domain-relevant features from a sparse set of visual tokens. Extensive experiments on 16 challenging remote sensing and medical imaging benchmarks demonstrate that ViTP establishes new state-of-the-art performance across a diverse range of downstream tasks. The code is available at https://github.com/zcablii/ViTP
@@ 1 Introduction
Decades of research in cognitive neuroscience have established that human visual perception is not a one-way street. It is a dynamic process arising from the constant interaction of two fundamental pathways: a bottom-up, data-driven pathway and a top-down, knowledge-driven one. The bottom-up process, famously elucidated by Hubel and Wiesel [62], involves a hierarchical construction of understanding: simple features like edges and orientations are extracted by early visual cortices and progressively assembled into complex object representations in higher cortical areas [117, 9]. This paradigm has been the primary inspiration for modern computer vision. For instance, prevailing Vision Transformer pretraining paradigms include supervised classification, MIM [55, 7, 132], contrastive learning [56, 127, 49], and image-text contrastive learning [129, 102, 192]. Recent multimodal large models [3, 25] usually utilize a vision tower to provide vision elements for high-level understanding. These paradigms are all fundamentally “bottom-up”: they operate on the premise that robust low-level perception is a prerequisite for high-level understanding [108, 97].
@@
Figure 1: (a) The synergistic relationship between perception and understanding in modern CV. Our proposed ViTP forges a previously underexplored link from high-level understanding to low-level perception. (b) Self-attention activation maps for the query patch (marked with a red cross). ViTP identifies fine-grained object parts that are high-level semantically related.
However, this bottom-up hierarchy only tells half the story. A wealth of research in cognitive neuroscience has demonstrated the profound top-down influence of high-level understanding, attention, and prior knowledge on foundational perception. Cognitive processes, guided by expectation and goals, actively modulate and refine how low-level visual information is processed, even at the earliest stages in the visual cortex [70, 130, 123]. This feedback mechanism allows the brain to focus on relevant details, fill in missing information, and interpret ambiguous stimuli with remarkable efficiency. While this reciprocal pathway—where understanding enhances perception—is integral to human vision, it remains critically underexplored in the computer vision models. This raises a compelling question: can abstract, high-level understanding be harnessed to directly guide a perception model toward learning more precise and domain-relevant representations?
@@
Figure 2: Comparison of pretraining paradigms for Vision Transformer (ViT) foundation models. ViTP employs an instruction-following objective to directly instil domain-specific perception capabilities into the vision backbone.
@@
Answering this question is particularly critical for pretraining powerful foundation models in specialized domains like remote sensing and medical imaging, where prevailing paradigms face formidable challenges. The supervised [152] pretraining risks overfitting to narrow label distributions; the MIM [55] may inadvertently neglect the fine-grained details of small but critical objects; image-image contrastive methods, while effective, are often notoriously difficult to optimize and demand substantial computational resources. Prevailing pretraining paradigms of image-to-text contrastive learning [129] or multimodal autoregressive Vision Transformer pretraining [44] represent steps toward reasoning-guided pretraining. However, these methods primarily aim for general-purpose image-text alignment rather than optimizing a vision backbone for specialized, fine-grained downstream tasks. Consequently, the global, image-level features learned by models like CLIP [129] or AIM [44] often prove suboptimal for dense prediction tasks such as semantic segmentation or object detection [89]. This limitation underscores that merely aligning global image and text representations is an insufficient strategy for forging a powerful, domain-specific perception backbone.
To bridge this gap, we introduce Visual insTruction Pretraining (ViTP), a novel, top-down pretraining paradigm that directly integrates high-level, instruction-based understanding into the perceptual feature learning process of a Vision Transformer backbone. As illustrated in Figure 2, ViTP diverges from prior methods by embedding the Vision Transformer within a larger Vision-Language Model (VLM) and pretraining it via a visual instruction-following objective. We construct our training data with image-text pairs derived directly from the target downstream tasks. During training, image tokens from the Vision Transformer and text tokens from a user query are fed jointly into a Large Language Model (LLM) [103]. The LLM’s generated response serves as the supervisory signal, compelling the Vision Transformer backbone to learn the complex data distributions of the downstream domain in an end-to-end fashion. This process endows the Vision Transformer with highly relevant, domain-specific perceptual capabilities. Furthermore, we introduce Visual Robustness Learning (VRL) by randomly dropping a large fraction of the Vision Transformer’s output image tokens before they are passed to the LLM. This constraint implicitly forces the Vision Transformer’s attention mechanism to encode more comprehensive and robust information within each of the remaining tokens, thereby enhancing the robustness and semantic richness of the learned visual features.
@@
  • We introduce ViTP, the first top-down pretraining paradigm that leverages understanding tasks to imbue a Vision Transformer backbone with high-level semantic perception , as illustrated in Figure 1.
  • We propose Visual Robustness Learning, a regularization method that encourages the Vision Transformer to learn more comprehensive and robust feature representations by operating on a sparse set of visual tokens.
@@ 2.1 Perception Foundation Model
Early approaches predominantly relied on supervised pretraining over large-scale labeled datasets such as ImageNet [58, 40]. While effective in learning semantically meaningful representations, these methods risk overfitting to the label space of the pretraining task, thereby limiting generalization to diverse downstream applications. Consequently, the field has increasingly shifted toward unsupervised learning paradigms that leverage unlabeled data. Among these, contrastive learning and masked image modeling (MIM) emerged as two dominant strategies. Contrastive learning methods, including MoCo [56], BYOL [47], and DINOv2 [127], learn discriminative representations by minimizing the distance between positive pairs (e.g., augmented views of the same image) while maximizing the distance between negative pairs in the embedding space. These methods emphasize semantic invariance and shown strong generalization capabilities across tasks. On the other hand, MIM-based approaches such as MAE [55], SimMIM [170] and UM-MAE [88] adopt a generative pretext task: a large portion of the input image is masked, and the model is trained to reconstruct the missing content. This encourages the model to learn holistic, contextually rich visual features by capturing spatial dependencies and fine-grained details. While both paradigms achieved impressive results on natural images, their direct transfer to specialized domains such as remote sensing or medical imagery often results in suboptimal performance [32, 102, 50]. This underscores the necessity for domain-specific adaptations in pretraining strategies.
@@ 2.2 Visual Instruction Tuning
The rise of powerful VLMs such as Gemini [147], InternVL [25], and Qwen-VL [4] are largely driven by visual instruction tuning [103]. This paradigm aims to align visual representations with a pretrained Large Language Model by projecting image features into the LLM’s embedding space. The model is then finetuned on a corpus of image-text instruction pairs, enabling the LLM to interpret visual inputs and respond to textual instructions. A key challenge in this paradigm is the scarcity of high-quality, large-scale visual instruction data. Foundational models like LLaVA [103] addressed this by creating synthetic datasets: they used a powerful language model (e.g., GPT-4 [1]) to generate diverse and complex reasoning-based datasets. In the visual instruction tuning standard setup, the vision encoder typically remains frozen, serving solely as a feature extractor, while all learning is concentrated in the LLM and the projection module. While this paradigm effectively leverages perception to enhance understanding, the reverse direction, i.e. using understanding to improve perception, remains largely underexplored. Our proposed ViTP addresses this gap by inverting the traditional visual instruction tuning pipeline. Instead of using images to tune an LLM, we utilize an LLM to guide the pretraining of the vision encoder.
More related work is provided in the Supplementary Material.
@@ 3 Method
Figure 4: A conceptual illustration of the ViTP framework. A Vision Transformer (ViT) backbone is embedded within a large VLM and then pretrained with domain-specific instruction following objective and Visual Robustness Learning (VRL). This process instils high-level semantic understanding into the ViT. The resulting weights are then used to initialize models for various downstream perception tasks.
In this section, we present our visual instruction pretraining (ViTP) framework, a novel paradigm designed to pretrain a ViT backbone. As illustrated in Figure 2, ViTP fundamentally differs from traditional pretraining by leveraging the understanding capabilities of modern VLMs. The entire process is driven by a “visual instruction following objective”, where the ViT learns to extract features that help an LLM answer questions about an image. To tailor the model for specific downstream applications, we outline a “data recipe” for curating a domain-specific pretraining dataset. During pretraining, we employ “Visual Robustness Learning (VRL)”, a regularization technique that drops image tokens to enhance the semantic robustness of the learned representations. The final pretrained ViT serves as a powerful foundation model, readily adaptable to a variety of downstream tasks.
@@ 3.1 Visual Instruction Following Objective
The central hypothesis of ViTP is that a ViT can learn relevant and expressive features if its training is guided by high-level understanding. To achieve this, we frame the pretraining as a visual instruction-following task. The framework, depicted in Figure 4, processes a domain-specific image through a ViT encoder to produce a sequence of image tokens. These tokens are projected into the LLM’s embedding space and concatenated with the tokenized text of an instruction. A Large Language Model (LLM) then processes this combined sequence to generate a response. The entire model is trained end-to-end, allowing the supervisory signal from the LLM’s response to directly optimize the ViT’s feature extraction process.
Our approach employs a continual pretraining strategy [51], starting with a well-trained, general-purpose VLM. We then continue its training on our curated domain-specific datasets. Such a continual pretraining strategy leverages the VLM’s vast pre-existing knowledge of general visual and linguistic patterns, providing a robust initialization. It significantly accelerates convergence during domain-specific pretraining, enhancing computational efficiency. Let the raw dataset be $\mathcal{D}_{\text{raw}}=\{(I,Q,R)\}$, where $I$ is an image, $Q$ is a text query (instruction), and $R$ is the ground-truth text response. This is processed into the final training set $\mathcal{D}=\{(x_{i},x_{t},y^{\star})\}$, where $x_{i}$, $x_{t}$, and $y^{\star}$ represent the processed image tokens, text tokens, and target response tokens, respectively.
@@ 3.1.3 LLM-based Supervision
The combined sequence $S_{llm}$ is processed by the LLM, which acts as a understanding engine to interpret the visual information from $x_{i}$ in the context of the instruction $x_{t}$. The model then auto-regressively generates an output sequence $O$. The entire system is optimized by minimizing the discrepancy between the generated output $O$ and the ground-truth target response $y^{\star}$. For instance, if the instruction asks to identify an object, $y^{\star}$ could be a textual description including its location in a structured format. During pretraining, the gradients from the output loss propagate back through the entire model, including the projection layer and the ViT backbone. We allow the weights of the ViT, the projection layer, and the LLM to be trainable. The optimization follows a standard supervised finetuning (SFT) objective, which minimizes the negative log-likelihood of the target sequence:
@@ 3.3 Pretraining Dataset Recipe
    Task Capability Alignment: The instruction-following tasks in the pretraining data should foster abilities required downstream. For instance, for downstream object detection tasks, including visual grounding and fine-grained VQA during pretraining empowers the backbone with localization and spatial understanding capabilities.
@@ 3.4 Downstream Finetuning
Once ViTP is complete, the pretrained ViT backbone is then extracted to serve as a powerful backbone network. For downstream applications, the pretrained ViT is integrated into a standard ViT-Adapter [23] and combined with task-specific heads. The entire model is then finetuned on the target downstream dataset. This transfer learning approach leverages the rich, instruction-aware representations learned during ViTP, enabling faster adaptation and superior performance on specialized tasks.